12.70 L of a gas has pressure of 0.63 atm. What is the volume of the gas at 105kPa
1
Expert's answer
2020-04-03T11:49:45-0400

p1xV1=p2xV2, 1atm=101.325 kpa.

p2=105/101.325 = 1.03626943 atm.

V2= p1xV1/p2=0.63 x 12.7 / 1.03626943 = 7.720965 L.